OpenStudy (anonymous):

Variable y varies directly with x2, and y = 96 when x = 4. What is the value of y when x = 2? y =

OpenStudy (nurali):

y=kx^2

OpenStudy (anonymous):

\[y \; \; \; \alpha \; \; \; x^2\]

OpenStudy (nurali):

96=kx^2

OpenStudy (anonymous):

So let the constant of proportionality is k.. \( y = kx^2\)

OpenStudy (nurali):

96=k(4)^2

OpenStudy (nurali):

96=k(16)

OpenStudy (nurali):

k=6

OpenStudy (anonymous):

k=6?

OpenStudy (nurali):

y=6x^2

OpenStudy (anonymous):

what does that simplify to

OpenStudy (anonymous):

thanks i was confused can you help in about 30-60 seconds

OpenStudy (nurali):

y=6(2)^2

OpenStudy (nurali):

y=6*4

OpenStudy (nurali):

y=24

OpenStudy (nurali):

So the answer to your problem is that when x = 2, y equals 24.

OpenStudy (nurali):

Hope this helps you to understand the problem.